a rectangle has a perimeter of 101 cm. the width is 0.01 of the length. what are the dimensions of the rectangle ?

Answer by tutorcecilia(2152)

P=2(lenght + width)
101cm = 2(x + .01x)
101=2(1.01x)
101=2.02x
50=x
.
.
Checking the answer:
x=50
P=2(50 + [(.01)(50)]
P=2(50+.5)
P=2(50.5)
P=101